A certain Call option and Put option for Walker Industries stock both have an exercise (strike) price of $35.00. The Call premium (price) is $3.21 and the Put premium (price) is $5.32. Assume the stock pays NO dividends, and that the risk-free rate is 4%. Both options expire in 41 days. 1. Using the put/call parity model, calculate the current stock price (S). (Show all work. Highlight in bold your answer.) [4 pts.] 2. Based upon your answer above for the stock price, which option (the Put or the Call) is in-the-money? Briefly explain your answer.
Answer #1

Using Put Call Parity
S+P=C+X/(1+r)^t
where
S is current stock price
P is put premium
C is call premium
X is exercise or strike price
t is time to expiry

S=C+X/(1+r)^t-P=3.21+35/(1+4%)^(41/365)-5.32=32.73614266


Put is in the money when strike price is more than stock price Hence, put is in the money